March BG Com Night 4, March 10 2009
Observers: J. Hill, A. Rakich (Tucson) and J. Storm (Potsdam)
Lucifer support: Nancy Ageorges (on-call)
Telescope Operator: A. Ceranski
SW Support: N. Cushing, C. Biddick (Tucson)
Telescope Support: J. Little
Mountain enshouded by fog, so it doesn't look like a good night for observing.
Still some ice on the dome. We entertained ourselves for a while with closed dome tests.
Details (Times reported in MST below):
Closed Dome Tests
21:00 - D. Ashby is rebalancing telescope after DX M2 swing arm was moved out.
TCS Test 7
20:30 Verified that RotateZ
command works correctly with today's patches. Positive RotateZ
moves M3's focus toward the front of the telescope.
21:10 And conducted some operator training about the idiosyncracies of OSS and M3 UMAC.
Test GCS Acquire Mode
21:17 pointto BS9122 ACQUIRE gs=0, acquire image is now displayed successfully.
21:20 pointto BS9122 ACTIVE gs=0, star verification failed, but it incorrectly found star centroids with the dome lights on.
21:23 pointto BS9122 ACTIVE gs=0, correctly did not find star in acquisition image with dome lights off.
TCS Tests 10 and 11
21:25 Aaron uses presetGUI to go to 8:00:00 +00:00:00.
21:30 offset 5 5 coord="DETXY" - fails because guiding not active.
21:36 Put GCS in starOverlay mode by editting lbt.conf. Stop and restart GCS.
21:37 pointto BS9122 GUIDE
Note that IIF offset indicator was still Red "failed" after the successful preset. Chris patched this later in the night.
22:08:40 offset 5 5 coord="DETXY"
which did: ircOffsetPointing 0. 5. 5. DETXY MOUNT true ABS left
Preset Coords 07:30:34.500 +29:51:12.00
IIF Achieved 07:30:35.214 +29:51:01.40
So the conclusion is that Test 10 is passed, but Test 11 has failed because the achieved coords failed when they should not have. Called Michele, and she will look at this tomorrow.
End of the closed dome testing
LUCIFER Software Tests
01:54 Marcus sends a successful preset TRACK from LUCIFER.
01:56 Marcus sends offset commands.
- 02 Mar 2009